Catarina and Tom want to buy a rug for a room that is 10 by 19 feet. They want to leave an even strip of flooring uncovered around the edges of the room.
How wide of a strip will they have if they buy a rug with an area of 52 square feet?
(10-2x)(19-2x)=52 <====not sure if that's how i set it up

Answer by scott8148(6628)

your setup is correct







-------- SOLUTION BELOW -------------








190 - 58x + 4x^2 = 52

2x^2 - 29x + 69 = 0

(2x - 23) (x - 3) = 0

x = 23/2 ___ not realistic

x = 3